OpenStudy (anonymous):

how do you find the cotangent of 5pi/4 on the unit circle

OpenStudy (campbell_st):

find the tangent of 5pi/4 and then take the reciprocal... its a 3rd quadrant angle...

OpenStudy (campbell_st):

since tan = opp/adj then cot = adj/opp

OpenStudy (campbell_st):

so you are basically finding the and pi/4 and th e answer is 1 since tan(pi/4) = 1, the cotangent is 1

OpenStudy (anonymous):

Find the exact value of cot 240° using the unit circle. Rationalize the denominator, if necessary

OpenStudy (campbell_st):

cot(240) = cot(60) so find tan(60) then take the reciprocal.